# CrossDehaze: Scaling Up Image Dehazing with Cross-Data Vision Alignment and Augmentation

### v1

@@ III-B Feature Extractor
[…]
@@
III-C Internal-Augmentor with Self-supervised Learning
For the internal dataset, we propose an internal data augmentation strategy to improve the model’s ability to preserve details and contrast in dehazed images. Building upon external dataset augmentation, it emphasizes attention to local details in images, making the model more robust. Specifically, we applied different degrees of data augmentation to the output results $\widehat{y}$ of the augmentation model. This is represented as follows:
@@ III-C Internal-Augmentor with Self-supervised Learning
Where $I_{aggr}$ is the image generated through strong augmentation based on $I_{weak}$, and $P_{aggr}$ represents the strong augmentation operation. Specifically, weak augmentation refers to random cropping, and strong augmentation $P_{aggr}$ is Gaussian blurring. Finally, we compute the L2 norm between $I_{weak}$ and $I_{aggr}$, obtained through weak-strong augmentation to optimize the model’s internal loss as:
@@
$\alpha=\alpha\times\frac{1}{2}\left(\cos\left(\pi\frac{s}{S}\right)+1\right),$ (8)
@@
where $\alpha$ is a hyperparameter that varies with cosine decay. In experiments, we set the initial value of $\alpha=0.1$, $s$ is the current training iteration and $S$ is the total number of training iterations. By utilizing the self-supervised learning with $\mathcal{L}_{internal}$, we can obtain the internal similarity between data, which allows the model to learn richer features and be more robust.
As shown in Fig. 5, we visualize the difference in image quality with and without internal data augmentation. The first and last columns represent the hazy images and ground truth images, respectively. The second column shows the dehazed images obtained using the external data augmentation method. The third column shows dehazed images obtained using both external data augmentor and the strong-weak augmentation self-supervised learning (SW-SSL) approach for internal data augmentor. As shown in Fig. 5, by using the internal data augmentation self-supervised method, the details of the out-of-distribution image are closer to the real haze-free image.
Figure 5: Dehazed images obtained with and without internal data augmentor.The second column shows the dehazed images obtained only after applying RSID Channel Transfomer (RSCT). The third column shows the dehazed images obtained after applying both RSID Channel Transfomer and Strong-Weak Self-supervised Learning (RSCT+SW-SSL). It can be seen that the results obtained with internal augmentation have more prominent details, closely resembling the real haze-free image (GT).
In Algorithm 2, we present the implementation of strong-weak self-supervised augmentation on internal data.
III-D Loss Function
The original dehazing loss compares the generated dehazed images with the ground truth haze-free images using L1 loss. Then, we incorporate the self-supervised loss obtained from the internal dataset into the original loss. Therefore, we obtain the total loss from both external and internal data as follows:
$\mathcal{L}_{total}=\mathcal{L}_{external}+\mathcal{L}_{internal}$ (9)
IV Experiment
IV-A Experimental Settings
Our model is implemented using PyTorch 1.13.0 on an NVIDIA RTX 3090 GPU. The model is trained using the AdamW optimizer [49] with a cosine annealing strategy [50]. The batch size and training epochs are set to 8 and 3000, respectively. The learning rate is reduced from an initial rate of 1e-4 to 1e-6. Additionally, we apply the aforementioned external augmentation method to the RSID dataset. Subsequently, we incorporate the revised RSID dataset as an augmented dataset into the NID dataset for training.
IV-B Dataset
For dataset comparison, we select two datasets:
The NID [2] consists of synthetic natural hazy images. It relies on the estimation of depth information. Therefore, the NID dataset first uses a depth prediction algorithm to estimate the depth maps of natural images. Then, natural hazy images are synthesized based on a haze simulation strategy. The NID consists of 1000 pairs of hazy-real image pairs.
[…]
For the NID, we select the first 900 images in the order of arrangement as the training set, while the remaining 100 images are used as the test set (N100). Therefore, for the RSCT 3:1 dataset, a total of 1200 images are used as the training set. For the RSCT 1:1 dataset, a total of 1800 images are used as the training set.
IV-C Evaluation Metrics and Comparison Objects
To evaluate our method’s performance, we adopt Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) as evaluation metrics. These metrics are commonly used as standards for evaluating image quality in image dehazing tasks.
PSNR(Peak Signal to Noise Ratio)is an objective measure for evaluating image quality. A higher value indicates better image quality, and it’s commonly used to quantify the visual difference between two images. Here, it represents the difference between the resulting dehazed image and the ground truth haze-free image.
$PSNR=10\times log_{10}(\frac{MAX^{2}}{MSE}),$ (10)
where MAX is the possible maximum pixel value in the image. MSE stands for Mean Squared Error, which represents the average of the squared differences between the pixel values of the true clear image and the dehazed image.
SSIM (Structural Similarity Index) is another metric used to measure the quality of images. The SSIM metric considers three aspects of similarity in images: luminance, contrast, and structure. Its calculation formula is as follows:
$\text{SSIM}(a,b)=\frac{(2\mu_{a}\mu_{b}+c_{1})(2\sigma_{ab}+c_{2})}{(\mu_{a}^{2}+\mu_{b}^{2}+c_{1})(\sigma_{a}^{2}+\sigma_{b}^{2}+c_{2})},$ (11)
@@ IV-C Evaluation Metrics and Comparison Objects
Figure 6: Dehazed Images Obtained by Different Methods on the NID
@@
Figure 7: Dehazed Images Obtained by Different Methods on the RSID
@@
Figure 8: Dehazed Images Obtained by Different Methods on the RSID
[…]
IV-D Comparison with Baseline Methods
We compare our method with state-of-the-art dehazing works. The experiments show that our results are significantly better than other methods, as shown in Tab. I with specific experimental results.
TABLE I: Quantitative Results on the NID
Method $PSNR\uparrow$ $SSIM\uparrow$
DEFADE [11] 15.80 0.8383
GDCP [6] 15.99 0.7433
zero-store [51] 16.09 0.7849
ROP [9] 16.39 0.8087
TCN [52] 16.36 0.7112
Haze-Lines [53] 17.96 0.8378
ALC [8] 19.03 0.8499
ZID [54] 19.62 0.8270
DeHamer [1] 23.11 0.9369
UHD [12] 23.72 0.9182
Trinity-Net [2] 28.80 0.9743
Ours 31.14 0.9896
TABLE II: Quantitative Results on the RSID
Method $PSNR\uparrow$ $SSIM\uparrow$
IDeRs [55] 13.60 0.6439
TCN [52] 14.21 0.6058
GRS-HTM [56] 14.80 0.5190
ROP [9] 15.57 0.7500
EVPM [57] 15.58 0.6889
SDCP [58] 16.05 0.6913
STD [59] 16.26 0.5592
zero-store [51] 16.65 0.7173
ZID [54] 18.99 0.7267
UHD [12] 22.06 0.8891
DeHamer [1] 23.61 0.8990
Trinity-Net [2] 23.78 0.9224
Ours 25.44 0.9399
@@ IV-D Comparison with Baseline Methods
TABLE III: Ablation Study Using Only Internal Data Augmentation and Self-Supervised Module
Internal-augmentor SW-SSL N100
@@
TABLE IV: Ablation Study Using Only External Data Augmentation
@@
TABLE V: Ablation study using both external and internal data augmentation methods
@@
SW-SSL RSCT3:1 RSCT1:1 $PSNR\uparrow$ $SSIM\uparrow$
@@
Observing Tab. I, we can see that our method first achieves a PSNR metric reaching 31.14. Moreover, it significantly outperforms other methods in terms of both PSNR and SSIM metrics. We observe that our method achieves a PSNR improvement of 2.34 dB compared to the Trinity-Net [2]. After our method, its SSIM value has increased by 0.0153 compared to Trinity-Net’s 0.9743. Compared to the UHD [12], our method has improved by 7.42 dB in terms of PSNR and 0.0714 in terms of SSIM. These results indicate that our method is better at restoring image visibility.
@@
As shown in Fig. 6, to specifically demonstrate the superiority of our method, we compared our model results with other methods. The first column represents the hazy image. The second, third, fourth, and fifth columns respectively depict the dehazed images obtained by the Zero-store [51], UHD [12], Dehamer [1], and Trinity-Net [2] methods. The sixth column displays the dehazed image achieved through our method. The last column shows the ground truth haze-free image. We compare the dehazed images obtained by our method with those obtained by other methods on the N100 test set. For example, from the images obtained above, it can be seen that our method approaches the real image more closely in terms of texture details. From the zoomed-in images, it can be observed that our dehazed images are noticeably closer to the real haze-free images in terms of texture and color. Moreover, the edge details are clearer compared to the Trinity-Net [2] method.
Figure 9: Dehazed Images Obtained After Internal Data Augmentation and Self-Supervised Module.The second column shows the dehazed images obtained without applying Strong-Weak Self-supervised Learning (w/o SW-SSL).The third column shows the dehazed images obtained after applying Strong-Weak Self-supervised Learning (w/ SW-SSL).
Figure 10: Dehazed images obtained using external data augmentation method with ratios of 3:1 and 1:1 respectively.The second column shows the dehazed images obtained without applying RSCT (w/o RSCT).The third column shows the dehazed images obtained after applying RSCT with ratios of 3:1 (w/RSCT 3:1). The fourth column shows the dehazed images obtained after applying RSCT with ratios of 1:1 (w/RSCT 1:1).
Figure 11: Dehazed images obtained using both external and internal data augmentation methods.The second column shows the dehazed images obtained without applying RSCT and SW-SSL (w/o RSCT+SW-SSL). The third column shows the dehazed images obtained after applying RSCT with ratios of 3:1 and SW-SSL (w/RSCT3:1+SW-SSL). The fourth column shows the dehazed images obtained after applying RSCT with ratios of 1:1 and SW-SSL (w/RSCT1:1+SW-SSL).
@@
We also apply our method to the remote sensing dataset RSID. Fig. 7, Fig. 8 specifically illustrates the comparison between our method and other methods. From the figures, it can be seen that the dehazed images obtained by the zero-store method [51] still contain haze, while the dehazed images obtained by the UHD method [12] exhibit partial distortion in terms of color. In comparison to DeHamer [1], in the results shown in the fourth column, there is still a small amount of residual haze along the edges. In comparison to Trinity-Net [2], in the results shown in the fifth column, Trinity-Net [2] results exhibit differences in local color compared to the ground truth haze-free image. In our method, the haze has been removed, and there is no distortion in color. The dehazed images obtained through our method are closest to the real haze-free images.
@@
In Tab. II, we observe that our method outperforms other methods significantly in both PSNR and SSIM metrics. For the zero-store [51], UHD [12], DeHamer [1], and Trinity-Net methods [2], we obtained actual quantitative results through reproduction. The experimental results of other methods are obtained based on the Trinity-Net [2] paper. Compared to Trinity-Net [2], our method achieved an improvement of 1.66 dB in the quantitative metric PSNR and an improvement of 0.0175 in SSIM. Compared to Dehamer [1], our method achieved an improvement of 1.83 dB in the quantitative metric PSNR and an improvement of 0.0409 in SSIM. These quantitative results indicate that our method also produces dehazed images with higher visual quality on the remote sensing dataset RSID.
@@ IV-E Ablation Study
To demonstrate the effectiveness of our proposed method, we conduct ablation studies to analyze the effects of different modules, including external data augmentation, internal data augmentation, and self-supervised methods. We add different modules to the network, including: (1) Only using internal data augmentation and self-supervised method. (2) Only using external data augmentation method. (3) Using both external data augmentation method and internal data augmentation and self-supervised method simultaneously.
We use L1 loss as the initial image reconstruction loss (Eq. 4), while in the self-supervised method, we use $\mathcal{L}_{internal}$ loss. We use the remaining one hundred images from the NID dataset (NID100) as the test set. The summary of the metrics for these models is presented in the following table.
@@
In Tab. III, we only used internal data augmentation and a self-supervised module in our method. Compared to not adding modules, using internal data augmentation and self-supervised modules results in an improvement of 0.14 dB in PSNR values on NID100. In Fig. 9, partial dehazed images obtained after internal data augmentation and self-supervised modules are shown.
@@
In Tab. IV, we solely employ external data augmentation in the model. Specifically, we incorporate external data augmentation by including the augmented RSID dataset in the training set. The training images from the NID dataset and the RSID dataset are divided into a ratio of 3:1 and 1:1, respectively, for training. In Tab. IV, compared to the results obtained without adding any modules, the results obtained with a 1:1 ratio show improvements of 0.65dB and 0.0021 in PSNR and SSIM metrics, respectively. Compared to the 3:1 ratio, the 1:1 ratio shows improvements of 0.24 and 0.0014 in PSNR and SSIM metrics, respectively. To some extent, it indicates that appropriately augmenting the dataset can make the model more robust. Fig. 10 shows the experimental results of haze removal with external data augmentation at ratios of 3:1 and 1:1, respectively.
@@
Then, we train the network using both external and internal augmentation methods simultaneously, as shown in Tab. V. We observe that the highest PSNR value of 31.14dB is achieved when simultaneously using the weak-strong augmentation self-supervised method and dividing the training data of NID and RSID in a 1:1 ratio during the channel transfer module. Compared to the results obtained without adding any modules, the PSNR value on the test set has increased by 1.08dB, and the SSIM has increased by 0.0025. Compared to the 3:1 split of training data between NID and RSID, the PSNR on the test set obtained through the 1:1 split has increased by 0.44dB. As shown in Fig. 11, the single dehazed image obtained after applying both modules exhibits higher visual quality and is closer to the true haze-free image in terms of overall color.
V Conclusion

### v2

@@ III-B Feature Representation Extraction
Then, we use the L1 loss function to train the feature extractor, obtaining the loss for external data:
@@
Figure 5: Dehazed images obtained with and without internal data augmentor.
III-C Internal- Augmentation with Self-supervised Learning
For the internal dataset, we propose an internal- augmentation strategy to improve the model’s ability to preserve details and contrast in dehazed images. Building upon external dataset augmentation, it emphasizes attention to local details in images, making the model more robust. Specifically, we apply different degrees of data augmentation to the output results $\widehat{y}$ of the augmentation model. This is represented as follows:
@@ III-C Internal- Augmentation with Self-supervised Learning
where $I_{aggr}$ is the image generated through strong augmentation based on $I_{weak}$, and $P_{aggr}$ represents the strong augmentation operation. Specifically, weak augmentation refers to random cropping, and strong augmentation $P_{aggr}$ is Gaussian blurring. In the Gaussian algorithm, our parameters are set as follows: the kernel size is set to 9, and the standard deviation is set to (0.1, 5). Finally, we compute the L2 norm between $I_{weak}$ and $I_{aggr}$, obtained through weak-strong augmentation to optimize the model’s internal loss as:
@@
$\alpha=\alpha\times\frac{1}{2}\left(\cos\left(\pi\frac{s}{S}\right)+1\right)$ (8)
@@
where $\alpha$ is a hyperparameter that varies with cosine decay. In experiments, we set the initial value of $\alpha=0.1$, $s$ is the current training iteration and $S$ is the total number of training iterations. As the number of training iterations increases, the cosine decay function ensures that the internal loss decreases along a smooth curve. It avoids sudden changes that could cause instability, ensuring the model’s stability during the training process. By utilizing the self-supervised learning with $\mathcal{L}_{internal}$, we can obtain the internal similarity between data, which allows the model to learn richer features and be more robust.
@@
Figure 6: Dehazed images obtained by state-of-the-art methods on the NID. Zooming up for better view.
@@
Figure 7: Dehazed images obtained by state-of-the-art methods on the RSID. Zooming up for better view.
@@
Figure 8: Dehazed images obtained by state-of-the-art methods on the RSID. Zooming up for better view.
Figure 9: Dehazed Images Obtained After Internal- Augmentation and Self-Supervised Module.
Figure 10: Dehazed images obtained using external- augmentation method with ratios of 3:1 and 1:1 respectively.
Figure 11: Dehazed images obtained using both external and internal- augmentation methods.
Figure 12: The training loss curve of our method.
TABLE I: Quantitative results on the NID.
Method $PSNR\uparrow$ $SSIM\uparrow$ $MSE(\times 10^{3})\downarrow$ $FSIM\uparrow$
Zero-Store [31] 16.09 0.7849 1.9159 0.9550
UCL-Dehaze [32] 18.03 0.8830 2.3157 0.9349
DeHamer [10] 23.11 0.9369 0.3959 0.9738
UHD [33] 23.72 0.9182 0.2921 0.9700
Trinity-Net [13] 28.80 0.9743 0.2581 0.9906
DEA-Net [34] 28.04 0.9814 0.3713 0.9944
MixDehazeNet [35] 29.74 0.9880 0.2552 0.9957
Ours 31.14 0.9896 0.1973 0.9962
TABLE II: Quantitative results on the RSID.
Method $PSNR\uparrow$ $SSIM\uparrow$ $MSE(\times 10^{3})\downarrow$ $FSIM\uparrow$
Zero-Store [31] 16.65 0.7173 2.0185 0.8904
UCL-Dehaze [32] 17.59 0.8090 2.4816 0.8739
UHD [33] 22.06 0.8891 1.7286 0.9421
DeHamer [10] 23.61 0.8990 1.1223 0.9475
Trinity-Net [13] 23.78 0.9224 1.0805 0.9541
MixDehazeNet [35] 24.20 0.9238 0.9422 0.9591
DEA-Net [34] 24.79 0.9343 0.8787 0.9646
Ours 25.44 0.9399 0.7481 0.9665
TABLE III: Efficiency analysis.
Model UHD [33] Trinity [13] DEA-Net [34] Ours
Params (M) 34.54 20.24 7.79 2.51
GFLOPS 106 144 92 110
Runtime(hours) 11 12 34 40
$PSNR$ 23.72 28.80 28.04 31.14
$SSIM$ 0.9182 0.9743 0.9814 0.9896
@@
TABLE IV: Ablation study using internal-augmentation and self-supervised module.
Internal-augmentor WS-SSL N100
@@
TABLE V: Ablation study using only external-augmentation.
@@
TABLE VI: Ablation study using both external and internal-augmentation methods.
@@
WS-SSL RSCT3:1 RSCT1:1 $PSNR\uparrow$ $SSIM\uparrow$
@@
TABLE VII: Ablation study using our method toward Trinity-Net on the NID.
Method $PSNR\uparrow$ $SSIM\uparrow$
Trinity-Net+origion DA 28.80 0.9743
Trinity-Net+our method 29.12 0.9755
TABLE VIII: Ablation study using our method in DEA-Net on the RSID.
Method $PSNR\uparrow$ $SSIM\uparrow$
DEA-Net+origion DA 24.79 0.9343
DEA-Net+our method 25.28 0.9382
TABLE IX: Ablation study on our method and fixed label supervision.
Method $PSNR\uparrow$ $SSIM\uparrow$
Fixed Label Supervision 30.86 0.9894
Ours 31.14 0.9896
As shown in Fig. 5, we visualize the difference in image quality with and without internal- augmentation. The first and last columns represent the hazy images and ground truth images, respectively. The second column shows the dehazed images obtained using the external- augmentation method. The third column shows dehazed images obtained using both external data augmentor and the weak-to-strong augmentation self-supervised learning (WS-SSL) approach for internal data augmentor. As shown in Fig. 5, by using the internal- augmentation and self-supervised learning (SSL), the details of the out-of-distribution image are closer to the real haze-free image.
In Algorithm 2, we present the implementation of weak-to-strong self-supervised augmentation on internal data. We incorporate the self-supervised loss obtained from the internal dataset into the original loss. Therefore, we obtain the total loss from both external and internal data as follows:
$\mathcal{L}_{total}=\mathcal{L}_{external}+\mathcal{L}_{internal}$ (9)
IV Experiment
IV-A Experimental Settings
Our model is implemented using PyTorch 1.13.0 on an NVIDIA RTX 3090 GPU. The model is trained using the AdamW optimizer [36] with a cosine annealing strategy [37]. The batch size and training epochs are set to 8 and 3000, respectively. The learning rate is reduced from an initial rate of 1e-4 to 1e-6. As shown in Fig. 12, we present the loss curve of our method during the training process. It can be observed that the training curve converges when the epoch reaches 3000. Additionally, we apply the aforementioned external augmentation method to the RSID. Subsequently, we incorporate the revised RSID as an augmented dataset into the NID for training.
IV-B Dataset
For dataset comparison, we select two datasets: The NID [13], and The RSID [13]. For the NID, we select the first 900 images in the order of arrangement as the training set, while the remaining 100 images are used as the test set (N100). Therefore, for the RSCT 3:1 dataset, a total of 1200 images are used as the training set. For the RSCT 1:1 dataset, a total of 1800 images are used as the training set.
IV-C Evaluation Metrics and Comparison Objects
To evaluate the performance of our method, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [38], Mean Squared Error (MSE), and Feature Similarity Index (FSIM) [39] as evaluation metrics.
IV-D Comparison with Baseline Methods
We compare our method with state-of-the-art dehazing works. The experiments show that our results are significantly better than other methods, as shown in Table I with specific experimental results.
@@ IV-D Comparison with Baseline Methods
Observing Table I, we can see that our method first achieves a PSNR metric that reaches 31.14. In addition, our method significantly outperforms other approaches in all four evaluation metrics. We observe that compared to Trinity-Net [13], our method achieves a 2.34 dB improvement in PSNR, an increase of 0.0153 in SSIM and an improvement of 0.0056 in FSIM. Additionally, the MSE value after applying our method is reduced by 0.0608 compared to the Trinity-Net [13]’s MSE of 0.2581. Compared to DEA-Net [34], our method has improved by 3.1 dB in terms of PSNR and 0.0082 in terms of SSIM. Compared to UCL-Dehaze [32], our method demonstrates complete superiority across all four metrics. In contrast to MixDehazeNet [35], our method achieves an improvement of 1.4 dB in the PSNR metric and an enhancement of 0.006 in SSIM. These results indicate that our method is better at restoring image visibility.
@@
As shown in Fig. 6, to specifically demonstrate the superiority of our method, we compared our model results with other methods, i.e., Zero-store [31],UCL-Dehaze [32] UHD [33], Dehamer [10], Trinity-Net [13], DEA-Net [34] and MixDehazeNet [35] . For example, from the images obtained above, it can be seen that our method approaches the real image more closely in terms of texture details. From the zoomed-in images, it can be observed that our dehazed images are noticeably closer to the real haze-free images in terms of texture and color. Moreover, the edge details are clearer compared to Trinity-Net [13] and DEA-Net [34] methods.
@@
We also apply our method to the remote sensing dataset (i.e., RSID). Fig. 7, Fig. 8 specifically illustrates the comparison between our method and other methods. From the figures, it can be seen that the dehazed images obtained by Zero-Store [31] and UCL-Dehaze [32] still contain haze, while the dehazed images obtained by the UHD method [33] exhibit partial distortion in terms of color. In comparison to DeHamer [10], in the results shown in the fourth column, there is still a small amount of residual haze along the edges. The results of Trinity-Net [13],DEA-Net [34] and MixDehazeNet [35] exhibit differences in local color compared to the haze-free ground truth image. In our method, the haze has been removed, and there is no distortion in color. The dehazed images obtained through our method are closest to the real haze-free images.
@@
In Table II, we observe that our method also significantly outperforms other approaches in all four evaluation metrics. For the Zero-Store [31], UHD [33], DeHamer [10], Trinity-Net [13],UCL-Dehaze [32], MixDehazeNet [35] and DEA-Net [34] methods, we obtained actual quantitative results through reproduction. We utilize same experimental setup as reported in the Trinity-Net paper [13], ensuring consistency in the experimental conditions. Compared to Trinity-Net [13], our method achieved an improvement of 1.66 dB in the quantitative metric PSNR and an improvement of 0.0175 in SSIM. Compared to UCL-Dehaze [32], our method also shows comprehensive improvements across all four metrics. In comparison to MixDehazeNet [35], our method achieves an improvement of 1.24 dB in the PSNR metric and an enhancement of approximately 0.016 in SSIM. These quantitative results indicate that our method also produces dehazed images with higher visual quality on the remote sensing dataset RSID.
Cross-model evaluation. To further evaluate our algorithm, we apply our method to both Trinity-Net and DEA-Net to better demonstrate the generalizability across other models. As shown in Table VII and VIII, on the NID, by applying our method, the PSNR of the NID test set increases by 0.32 dB, and the SSIM improves by 0.0012. Compared to the original data augmentation method, using our method increases the PSNR of the RSID test set by approximately 0.5 dB and the SSIM by 0.0039. This demonstrates that our method is also applicable to other models and enhances the robustness of the models.
Comparison with the full supervised framework. To further verify the effectiveness of our self-supervised learning method, we compare it with the method using fixed ground truth $y$. As shown in Table IX, our method achieves a PSNR improvement of 0.28 dB and an SSIM improvement of 0.0002 on the NID test set compared to the supervised framework (using fixed $y$). The experimental results demonstrate that the dynamically generated $I_{week}$ and $I_{aggr}$ outperform the fixed label supervision in capturing local structural features. This mechanism encourages the network to learn robust representations from its own outputs, indicating that our method more effectively integrates the dynamic predictions of the model during training.
Efficiency analysis. As shown in Table III, we conduct an efficiency analysis of our method compared to other approaches. Our method is evaluated based on the NID (1:1) dataset. Although our method has higher runtime and computational complexity (GFLOPS), the number of parameters in our model is significantly lower than that of other methods. Compared to Trinity-Net [13] and DEA-Net [34], our method maintains the highest evaluation metrics while having fewer parameters than both. The experimental results confirm that our method offers high accuracy and practical usability.
@@ IV-E Ablation Study
To demonstrate the effectiveness of our proposed method, we conduct ablation studies to analyze the effects of each module.
@@
Effectiveness of internal-augmentation. In ablation study, we use L1 loss as the initial image reconstruction loss (Eq. 4), while in the self-supervised method, we use $\mathcal{L}_{internal}$ loss. The remaining one hundred images from the NID dataset (NID100) are used as the test set. In Table IV, we only use internal- augmentation and a self-supervised module in our method. Compared to not adding modules, using internal- augmentation and self-supervised modules results in an improvement of 0.14 dB in PSNR values on NID100. In Fig. 9, partial dehazed images obtained after internal- augmentation and self-supervised modules are shown.
@@
Effectiveness of external-augmentation. In Table V, we solely employ external- augmentation in the model. Specifically, we incorporate external- augmentation by including the augmented RSID dataset in the training set. The training images from the NID and the RSID are divided into a ratio of 3:1 and 1:1, respectively, for training. In Table V, compared to the results obtained without adding any modules, the results obtained with a 1:1 ratio show improvements of 0.65dB and 0.0021 in PSNR and SSIM metrics, respectively. Compared to the 3:1 ratio, the 1:1 ratio shows improvements of 0.24 and 0.0014 in PSNR and SSIM metrics, respectively. RSCT 3:1 only expanded 300 gamma-corrected images, resulting in a total of 1200 images, while RSCT 1:1 expanded 900 gamma-corrected images, resulting in a total of 1800 images. As a result, RSCT 3:1 has a smaller dataset size and less data diversity compared to RSCT 1:1. This leads to the model learning fewer features during training than RSCT 1:1, reducing the model’s robustness and consequently resulting in lower performance on the test set compared to the results obtained from training with RSCT 1:1. Additionally, Fig. 10 shows the experimental results of haze removal with external- augmentation at ratios of 3:1 and 1:1, respectively.
@@
Effectiveness of external- and internal- augmentation. Then, we train the network using both external and internal augmentation methods simultaneously, as shown in Table VI. We observe that the highest PSNR value of 31.14dB is achieved when simultaneously using the weak-strong augmentation self-supervised method and dividing the training data of the NID and the RSID in a 1:1 ratio during the channel transfer module. Compared to the results obtained without adding any modules, the PSNR value on the test set has increased by 1.08dB, and the SSIM has increased by 0.0025. Compared to the 3:1 split of training data between the NID and the RSID, the PSNR on the test set obtained through the 1:1 split has increased by 0.44dB. As shown in Fig. 11, the single dehazed image obtained after applying both modules exhibits higher visual quality and is closer to the true haze-free image in terms of overall color.
V Conclusion and Limitation